OpenStudy (anonymous):

Solve for x x^2(x-3)^2-4(x-3)^2=0

OpenStudy (anonymous):

\[x ^{2}(x-3)^{2}-4(x-3)^{2}=0\]

OpenStudy (anonymous):

I have \[x ^{2}(x-3)(x-3)-4(x-3)(x-3)=0\] but I don't know what to do next

OpenStudy (anonymous):

That wasn't really necessary :) It may not be obvious yet, but try letting \[\large u=(x-3)^2\] Then \[\large x^2(x-3)^2-4(x-3)^2=x^2u-4u\]

OpenStudy (anonymous):

And you can see that u can simply be factored out like this... \[\large u(x^2-4)\]

OpenStudy (anonymous):

I am really confused by what you just posted

OpenStudy (anonymous):

And then revert u back to its original meaning, sort of... \[\large (x-3)^2(x^2-4)\] What I did was factor out the entire expression (x-3)^2

OpenStudy (anonymous):

You don't have to go through the u-part but I did it so that it's easier to see that it can be factored out :)

OpenStudy (anonymous):

ok but from the step that I was at before do you then do x^2(x^2-6x-6)-4(x^2-6x-6)=0

OpenStudy (anonymous):

:) First of all (x-3)^2 = x^2 - 6x + 9 Second, you can certainly do that, but you'll have to factor that out again later, so why bother? :)